kevin has $9,300 to invest. he decides to invest a portion of ther $9,300 at 4.7% for 1 year and the rest at 5.9% each for 1 year. if the total interest amount that kevin earns is $507.90, how much was invested at 4.7% and how much was invested at 5.9%?
pleaseeeeeeeeee helpppppppppp

Answer by funmath(2933)

Let the principle you invest at 4.7%=x
The rest you invest at 5.9%=9,300-x
time=1
Interest=Principle*rate*time
Interest at 4.7%=(x)*(.047)*(1)=.047x
Interest at 5.9%=(9,300-x)*(.059)*(1)=.059(9,300-X)
Total Interest: 507.90
So your problem to solve is:
.047x+.059(9,300-x)=507.90
.047x+.059(9,300)+.059(-x)=507.90
.047x+548.7-.059x=507.90
(.047-.059)x+548.7=507.90
-.012x+548.7=507.90
-.012x+548.7-548.7=507.90-548.7
-.012x=-40.8
-.012x/-.012=-40.8/-.012
x=3,400
The amount invested at 4.7%=$3,400
The amount invested at 5.9%=9,300-3,400=$5,900
If you truely hate decimals, you can eliminate your decimals for your first step by multiplying both sides of the equation by 1000 and then solving.
Happy investing!